# Review of Degenerate Higher Order Scalar Tensor Theories in Cosmology

### v1

@@ 6 Alternatives to inflation
$S=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+\mathcal{L}_{H2}+\mathcal{L}_{H3}+\mathcal{L}_{D}\right)~{},$ (6.1)
where $\mathcal{L}_{H2}$ and $\mathcal{L}_{H3}$ and $\mathcal{L}_{D}$ are given by
$\mathcal{L}_{H2}$ $=K(\pi,X)=-g_{1}(\pi)X+g_{2}(\pi)X^{2}~{},$ (6.2)
$\mathcal{L}_{H3}$ $=G(X)\Box X~{},~{}G(X)=\gamma X^{2}~{},$ (6.3)
$\mathcal{L}_{D}$ $=\frac{R}{2}h-\frac{h}{4X}\left[\pi_{\mu\nu}\pi^{\mu\nu}-(\Box\pi)^{2}\right]+\frac{h-2Xh_{X}}{4X^{2}}\left[\pi_{\mu}\pi^{\mu\rho}\pi_{\rho\nu}\pi^{\nu}-(\Box\pi)\pi^{\mu}\pi_{\mu\nu}\pi^{\nu}\right]\,,$ (6.4)
@@
$g_{1}(\pi)=\frac{3}{2}f^{2}e^{4\pi}\frac{1+e^{2\pi}}{1+e^{m\pi}}~{},~{}g_{2}(\pi)=e^{2\pi}\frac{1+e^{2\pi}}{1+e^{4\pi}}~{},$ (6.5)
@@
The parameter $m>6$ determines the behaviour when the universe exiting the quasi-Minkowskian state and $m=7$ through is used in the analysis.
For the background, the Friedmann equations are obtained
$3H^{2}=\rho_{\pi}~{},~{}-2\dot{H}=\rho_{\pi}+p_{\pi}~{},$ (6.7)
where $H$ is the Hubble parameter, $\rho_{\pi}$ and $p_{\pi}$ are the density and pressure of the matter field $\pi$, expressed by
$\rho_{\pi}=-\frac{1}{2}g_{1}\dot{\pi}^{2}+\frac{3}{4}g_{2}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},$ (6.8)
$p_{\pi}=-\frac{1}{2}g_{1}\dot{\pi}^{2}+\frac{1}{4}g_{2}\dot{\pi}^{4}-\gamma\dot{\pi}^{4}\ddot{\pi}\,,$ (6.9)
independent of the DHOST term (6.4) [117, 118]. Assuming as the initial condition of scalar field $\pi$ at far past to be $\pi\ll-1$, the first Friedmann equation can be written as
$3H^{2}=-\frac{3}{4}f^{2}e^{4\pi}\dot{\pi}^{2}+\frac{3}{4}e^{2\pi}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},$ (6.10)
@@
$p_{\pi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)\,.$ (6.11)
@@
$-2\dot{H}\simeq p_{\pi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)~{},$ (6.12)
@@
$H=\frac{h_{0}}{(-t)^{5}}~{},~{}h_{0}\equiv\frac{1}{10}\left(\frac{1}{2f^{2}}+\gamma\right)~{}.$ (6.13)
@@
$\rho_{\pi}=\frac{3}{4}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},~{}p_{\pi}=\frac{1}{4}\dot{\pi}^{4}-\gamma\dot{\pi}^{4}\ddot{\pi}~{},$ (6.14)
@@
$a\propto t^{\frac{1}{2}}~{},~{}\dot{\pi}\propto t^{-\frac{1}{2}}\,.$ (6.15)
@@
The power spectrum for scalar and tensor perturbations of the model can be investigated by computing the second order action for linear perturbations, similarly to the standard inflationary scenario (Sec. 5) in terms of the comoving curvature perturbation $\zeta$ and then performing the field quantisation using the Mukhanov-Sasaki prescription,
@@
$\frac{z_{s}^{2}}{2a^{2}}$ $=3+2\frac{\dot{\pi}^{2}(-g_{1}+3g_{2}\dot{\pi}^{2})+18H\gamma\dot{\pi}^{5}-6H^{2}}{(\gamma\dot{\pi}^{5}-2H)^{2}}\,,$ (6.17)
$\left(-\frac{z_{s}^{2}}{2a^{2}}\right)c_{s}^{2}$ $=1+h+\frac{2}{a}\frac{d}{dt}\frac{a\left(h_{X}\dot{\pi}^{2}-h-1\right)}{2H-\gamma\dot{\pi}^{5}}~{}.$ (6.18)
In the pre-emergent period $t<0$, the universe is almost static and $d\eta=dt/a\simeq dt$ and
@@
The wavemodes of observational interests all become superhorizon in the pre-emergent period $t<0$ [118]. Hence, the dynamics of scalar perturbation have to be evaluated from past infinity to the critical time scale $\eta_{E}=t_{E}$, where the emergent solution breaks down. After $t=t_{E}$, the universe quickly turns to a radiation dominated period, and the scalar perturbation freezes.
The scalar power spectrum can be evaluated using Eq. (5.22). By using vacuum initial conditions it can be easily shown that it is scale invariant as required.
The tensor power spectrum can be computed analogously and the tensor spectral index is $n_{t}=\frac{d\ln P_{t}}{d\ln k}=2$, consistent with previous studies [125, 127].
@@ 7 Conclusions
In this review we have presented the DHOST theories – the most general scalar-tensor theories involving one scalar field, and propagating one scalar degree of freedom and two tensor ones, avoiding the generation of ghosts. These theories have been developed from the middle of the 2010s, after the observation that they do not need to involve only second-order Euler-Lagrange equations in order to satisfy this constraint, but that this can also happen when the functional coefficients in the Lagrangian satisfy certain degeneracy conditions. They include previously studied Horndeski and beyond-Horndeski theories, and while a large literature is available covering these less general models, we have not discussed their observational effects in this work, focusing only on the DHOST literature.
@@
The research on these theories has developed into two directions in recent years. Firstly, they have been used to study the late-universe and to parameterise dark energy, incorporating previously studied models, exhibiting compatibility with the current cosmological observations and to model astrophysical objects, such as Newtonian and neutron stars as well as black holes. Secondly, they have been applied to the early universe, in the context of its early-time accelerated expansion.
@@
This review covers both these aspects. We discuss the construction of the DHOST theories as generalisations of Horndeski and beyond-Horndeski theories, their classification, observational implications in the late universe, the use of these theories as models of dark energy and cosmological constraints, as well as inflationary model building.
@@ Acknowledgments
This work was supported by a United Kingdom Research and Innovation (UKRI) Future Leaders Fellowship [Grant No. MR/V021974/2].

### v3

@@ 7 Other Developments for the Early Universe
$S=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+\mathcal{L}_{H2}+\mathcal{L}_{H3}+\mathcal{L}_{D}\right)\,,$ (7.1)
where $\mathcal{L}_{H2}$, $\mathcal{L}_{H3}$ and $\mathcal{L}_{D}$ are given by
$\mathcal{L}_{H2}$ $=K(\phi,X)=-g_{1}(\phi)X+g_{2}(\phi)X^{2}\,,$ (7.2)
$\mathcal{L}_{H3}$ $=G(X)\Box X~{},~{}G(X)=\gamma X^{2}\,,$ (7.3)
$\mathcal{L}_{D}$ $=\frac{R}{2}h-\frac{h}{4X}\left[\phi_{\mu\nu}\phi^{\mu\nu}-(\Box\phi)^{2}\right]+\frac{h-2Xh_{X}}{4X^{2}}\left[\phi_{\mu}\phi^{\mu\rho}\phi_{\rho\nu}\phi^{\nu}-(\Box\phi)\phi^{\mu}\phi_{\mu\nu}\phi^{\nu}\right]\,,$ (7.4)
@@
$g_{1}(\phi)=\frac{3}{2}f^{2}e^{4\phi}\frac{1+e^{2\phi}}{1+e^{m\phi}}\,,~{}g_{2}(\phi)=e^{2\phi}\frac{1+e^{2\phi}}{1+e^{4\phi}}\,,$ (7.5)
@@
The parameter $m>6$ determines the behaviour when the universe is exiting the quasi-Minkowskian state. Throughout the analysis, $m=7$ through is used.
For the background, the Friedmann equations are obtained,
$3H^{2}=\rho_{\phi}~{},~{}-2\dot{H}=\rho_{\phi}+p_{\phi}\,,$ (7.7)
where $\rho_{\phi}$ and $p_{\phi}$ are the density and pressure of the matter field $\phi$, expressed by
$\rho_{\phi}=-\frac{1}{2}g_{1}\dot{\phi}^{2}+\frac{3}{4}g_{2}\dot{\phi}^{4}+3H\gamma\dot{\phi}^{5}\,,$ (7.8)
$p_{\phi}=-\frac{1}{2}g_{1}\dot{\phi}^{2}+\frac{1}{4}g_{2}\dot{\phi}^{4}-\gamma\dot{\phi}^{4}\ddot{\phi}\,,$ (7.9)
independent of the DHOST term (7.4) [165, 166]. Assuming as the initial condition of scalar field $\phi$ in the far past to be $\phi\ll-1$, the first Friedmann equation can be written as
$3H^{2}=-\frac{3}{4}f^{2}e^{4\phi}\dot{\phi}^{2}+\frac{3}{4}e^{2\phi}\dot{\phi}^{4}+3H\gamma\dot{\phi}^{5}\,,$ (7.10)
@@
$p_{\phi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)\,.$ (7.11)
@@
$-2\dot{H}\simeq p_{\phi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)\,,$ (7.12)
@@
$H=\frac{h_{0}}{(-t)^{5}}~{},~{}h_{0}\equiv\frac{1}{10}\left(\frac{1}{2f^{2}}+\gamma\right)\,.$ (7.13)
@@
$\rho_{\phi}=\frac{3}{4}\dot{\phi}^{4}+3H\gamma\dot{\phi}^{5}\,,~{}p_{\phi}=\frac{1}{4}\dot{\phi}^{4}-\gamma\dot{\phi}^{4}\ddot{\phi}\,,$ (7.14)
@@
$a\propto t^{\frac{1}{2}}~{},~{}\dot{\phi}\propto t^{-\frac{1}{2}}\,.$ (7.15)
@@
The power spectrum for scalar and tensor perturbations of the model can be investigated by computing the second order action for linear perturbations, similarly to the standard inflationary scenario (Sec. 6), in terms of the comoving curvature perturbation $\zeta$ and then performing the field quantisation using the Mukhanov-Sasaki prescription,
@@
$\frac{z_{s}^{2}}{2a^{2}}$ $=3+2\frac{\dot{\phi}^{2}(-g_{1}+3g_{2}\dot{\phi}^{2})+18H\gamma\dot{\phi}^{5}-6H^{2}}{(\gamma\dot{\phi}^{5}-2H)^{2}}\,,$ (7.17)
$\left(-\frac{z_{s}^{2}}{2a^{2}}\right)c_{s}^{2}$ $=1+h+\frac{2}{a}\frac{d}{dt}\frac{a\left(h_{X}\dot{\phi}^{2}-h-1\right)}{2H-\gamma\dot{\phi}^{5}}~{}.$ (7.18)
In the pre-emergent period $t<0$, the universe is almost static, $d\eta=dt/a\simeq dt$ and
@@
The wavemodes of observational interests all become superhorizon in the pre-emergent period $t<0$ [166]. Hence, the dynamics of scalar perturbation have to be evaluated from past infinity to the critical time scale $\eta_{E}=t_{E}$, where the emergent solution breaks down. After $t=t_{E}$, the universe quickly turns to a radiation dominated epoch, and the scalar perturbation freezes.
The scalar power spectrum can be evaluated using Eq. (6.24). By using the vacuum initial conditions it can be easily shown that it is scale invariant as required. The tensor power spectrum can be computed analogously and the tensor spectral index is $n_{t}=\frac{d\ln P_{t}}{d\ln k}=2$, consistent with previous studies [35, 174].
Another recent contribution regards the influence of higher order terms in the DHOST Lagrangian on non-perturbative solutions, which drastically change the nature of gravitational waves in the non-linear regime, with consequences for the early cosmology [175, 176].
Also, a Horndeski/DHOST phase in the early universe could enhance the production of secondary gravitational waves. This effect could be used to test scalar-tensor theories in the early universe [177].
@@ 8 Conclusions
In this review we have presented the DHOST theories – the most general scalar-tensor theories involving one scalar field and its derivatives up to second order, and propagating one scalar degree of freedom and two tensor ones, avoiding the generation of ghosts. They are generalisations of other previously studied theories, and in particular Horndeski theories.
Horndeski theories, first described in 1974, and rediscovered in the 2000s, have long been thought to be the most general scalar-tensor theories having this property because the requirement of having second-order Euler-Lagrange was believed necessary. The building of beyond-Horndeski theories and the understanding of the mechanisms involved in avoiding the propagation of ghosts through degeneracies between the parameters of the models have led the way to the development from the middle of the 2010s of DHOST theories, which include as special cases the previously mentioned ones. Thus, these theories provide a general and flexible framework for parameterising the universe. While a large literature is available covering the less general models, we have not discussed their observational effects in this work in detail, focusing primarily on the DHOST literature.
@@
In recent years, the research on these theories has developed into two directions. Firstly, they have been used to study the late universe and to parameterise dark energy, incorporating previously studied models, exhibiting compatibility with the current cosmological observations and to model astrophysical objects, such as Newtonian and neutron stars as well as black holes. Secondly, they have been applied to the early universe, in the context of its early-time accelerated expansion.
@@
This review covers both these aspects. We discussed the construction of the DHOST theories as generalisations of Horndeski and beyond-Horndeski theories, their classification, observational implications in the late universe, the use of these theories as models of dark energy and cosmological constraints, as well as inflationary model building.
Thus, research related to these theories has initially focused on the late universe, with the aim of understanding current cosmological acceleration, and of testing gravity. Although very general, in order to be compatible with current cosmological and astrophysical observations, the functions appearing in the DHOST theories have to satisfy certain constraints. In particular, the detection of primordial gravitational waves in the GW170817 event, has significantly restricted the parameter space of the models, from both the gravitational wave speed and the avoidance of gravitational waves decay into dark energy. There is still a debate whether these constraints are relevant for cosmological scales, due to the extrapolation of scales involved.
Another avenue of research in DHOST theories which has been active is related to astrophysical objects and in particular to black hole physics. The vicinity of black holes, where strong fields are present, is particularly suited to study deviations from GR, which could be investigated with gravitational wave physics. The different predictions in compact objects of DHOST theories compared to GR can also provide explanations for astrophyical events, such as the one measured by GW190814 event, which has a mass and a radius significantly larger than in GR. This could also be the case for the pulsar PSR J1614-2230. Departures from GR can potentially be measured in neutron stars.
DHOST theories can also be successfully used in the early universe. Models of inflation were built using such theories, and also alternatives to it. The inflationary models have been shown to satisfy the CMB constraints at the power spectrum and bispectrum levels. As the standard model of inflation has an initial Big Bang singularity, alternatives to it have been developed that avoid this issue, such as the bouncing universe scenario, where DHOST theories are well suited. This theories can have distinct phenomenology to standard inflation, and future probes will be able to distinguish between them.
Thus, DHOST theories could be very useful in exploring the fundamental nature of gravity, in fact the motivation for which models of modified gravity have been developed. Future cosmological and astrophysical probes will be able to increase our understanding of the nature of gravity.
@@ Acknowledgments
This work was supported by a United Kingdom Research and Innovation (UKRI) Future Leaders Fellowship [Grant No. MR/V021974/2]. The author thanks Philippe Brax for the introduction to the topic of DHOST theories and for working together in opening up new aspects of interest.
Data Access Statement
No data were created or analysed in this study.